OpenStudy (anonymous):

the product of (3+ square root of 5) and (3-square root of 5 )

OpenStudy (anonymous):

4

OpenStudy (anonymous):

how did you get tht

OpenStudy (anonymous):

yes, its factored as a difference of squares so you get a^2 which is 3 squared (9) minus root of five squared which is five so 9-5=4

OpenStudy (anonymous):

\[(3+\sqrt{5})(3-\sqrt{5})\]square the 3s and the root 5 9-5=4

OpenStudy (anonymous):

(3-V5)*(3+V5) it is in the form of (a-b)(a+b) = (a^2-b^2)